OpenStudy (cutestgirlaround):

Solve for x. −2/3(3x−4)+3x=5/6 A. X= - 19/6 B. X= - 11/6 C. X= - 21/6 D. X= - 29/6

Awolflover1 (awolflover1):

\[\frac{ -2 }{ 3 } (3x−4)+3x= \frac{ 5 }{ 6 }\]

OpenStudy (cutestgirlaround):

yeah that's what it is thanks!

OpenStudy (cutestgirlaround):

But what's the answer?

Awolflover1 (awolflover1):

\[( \frac{ -2 }{ 3 } )(3x)+( \frac{ -2 }{ 3 } )(−4)+3x= \frac{ 5 }{ 6 }\]

Awolflover1 (awolflover1):

Distribute^ \[−2x+ \frac{ 8 }{ 3 }+3x= 5 6\] \[(−2x+3x)+( \frac{ 8 }{ 3 } )= 5 6\] Combine Like Terms

Awolflover1 (awolflover1):

Subtract 8/3 from both sides. \[x+ \frac{ 8 }{ 3 } = 5 6\]

Awolflover1 (awolflover1):

\[x+ \frac{ 8 }{ 3 }− \frac{ 8 }{ 3 } = \frac{ 5 }{ 6 } −\frac{ 8 }{ 3}\]

Awolflover1 (awolflover1):

\[\frac{ 5 }{ 6} − \frac{ 8 }{ 3} = ? \]

OpenStudy (cutestgirlaround):

would it be 5/6 - 16/6 = 11/6?

Awolflover1 (awolflover1):

yes

OpenStudy (cutestgirlaround):

thanks

Awolflover1 (awolflover1):

but 11/6 is negative so -11/6

Awolflover1 (awolflover1):

yw =)